Jimstan & Jimstan Corp. can sell a new 10-year bond with an annual coupon of 5.4% and a face value of $1,000 for $1,127.9. The company will incur flotation costs of $40 per bond and has a tax rate of 28%.

(1) What are the net proceeds from selling the bond?

(2) What is the company's pre-tax cost of debt?

(3) What is the company's after-tax cost of debt?
Answer #1

1)

Net proceeds = $1,127.9 - $40

Net proceeds = $1,087.9

2)

Pre tax cost of debt = 4.30%

Coupon = 0.054 * 1000 = 54

Keys to use in a financial calculator:

FV 1000

PV -1,087.9

PMT 54

N 10

CPT I/Y

3)

After tax cost of debt = 0.043 (1 - 0.28)

After tax cost of debt = 0.0310 or 3.10%